Santa Fe Company was started on January 1, Year 1, when it acquired $9,200 cash by issuing common stock. During Year 1, the company earned cash revenues of $5,300, paid cash expenses of $3,350, and paid a cash dividend of $900. Based on this information,

Multiple Choice

  • The Year 1 statement of cash flows would show net cash inflow from financing activities of $8,300.

  • The Year 1 statement of cash flows would show a net cash flow from financing activities of $9,200.

  • The Year 1 income statement would show net income of $1,050.

  • The balance sheet at December 31, Year 1 would show total equity of $14,500.
Answer #1

Net cash flows from operating activities

= 5300 - 3350 = 1950

Net cash flows from Financing Activities

= 9,200 - 900 = 8,300

Option A is the answer